obtain the value of the constant c for which the function u(x,t) = cosαcxsinαt is a solution of the wave equation ∂ ^2u/∂t^2 = c^2∂ ^2u/∂ x^2
1
Expert's answer
2019-03-19T10:41:51-0400

Solution:

If

"u(x,t)=cos acx sin at(1)"

is a solution of the wave equation

"\frac{ ∂ ^2u}{∂t^2}=\frac{c^2∂ ^2u}{∂ x^2},(2)"

then


"\frac{ ∂ u}{∂t}=acos acx cos at,(3)"

"\frac{ ∂ ^2u}{∂t^2}=-a^2cos acx sin at,(4)"

"\frac{ ∂ u}{∂x}=-acsin acx sin at,(5)"

"\frac{ ∂ ^2u}{∂x^2}=-a^2c^2cos acx sin at,(6)"

and

"-a^2cos acx sin at=-a^2c^4cos acx sin at.(7)"

Then from equation (7) we find c:

"c^4=\frac{-a^2cos acx sin at}{-a^2cos acx sin at},"

"c^4=1,"

"c=1\\ {and }\\ c=-1."

Answer: c=1 and c=-1.